OpenStudy (anonymous):

-6p - 21 = 3p - 12 would i combine like terms for this one

OpenStudy (anonymous):

You would get your like terms on the same side first and then combine.

OpenStudy (anonymous):

okay my anwser is three but shouldnt it be neg three

OpenStudy (anonymous):

I got -1.

OpenStudy (anonymous):

-6p - 21 = 3p - 12 -6p-21+21=3p-12+21 -6p=3p+9 -6p-3p=3p-3p+9 -9p=9 -9p/-9=9/-9 p=-1

OpenStudy (anonymous):

wow i have been out of school way to long lol thank you im struggling with all this stuff

OpenStudy (anonymous):

No problem. :) Let me know if you need anymore help.

OpenStudy (anonymous):

Work-Out Corner has 5 more than 3 times as many exercise bicycles as The Gym. Together they have 21 bicycles. Solve the equation x + 3x + 5 = 21 to find the number of bicycles at Work-Out Corner.

OpenStudy (anonymous):

how do i work this one out

OpenStudy (anonymous):

x + 3x + 5 = 21 4x+5=21 4x+5-5=21-5 4x=16 4x/4=16/4 x=4

OpenStudy (anonymous):

12 + 0.35x = 20.05

OpenStudy (anonymous):

12 + 0.35x = 20.05 12-12+0.35x=20.05-12 0.35x=8.05 0.35x/0.35=8.05/0.35 x=23

OpenStudy (anonymous):

0.6(y + 3) = 4.8

OpenStudy (anonymous):

i got this wrong too

OpenStudy (anonymous):

0.6(y + 3) = 4.8 0.6y+1.8=4.8 0.6y+1.8-1.8=4.8-1.8 0.6y=3 0.6y/0.6=3/0.6 y=5

OpenStudy (anonymous):

6 = 2(x + 8) - 5x

OpenStudy (anonymous):

6 = 2(x + 8) - 5x 6=2x+16-5x -16+6=-3x+16-16 -10=-3x -10/-3=-3x/-3 10/3=x

OpenStudy (anonymous):

9(e + 5) = 225

OpenStudy (anonymous):

9(e + 5) = 225 9e+45=225 9e+45-45=225-45 9e=180 9e/9=180/9 e=20

OpenStudy (anonymous):

- 1/3m – 7 = 5 last one i swear

OpenStudy (anonymous):

- 1/3m – 7 = 5 -1/3m-7+7=5+7 -1/3m=12 -1/3m/-1/3=12/-1/3 m=-36